# SGFormer: Spherical Geometry Transformer for 360 Depth Estimation

### v2

@@ I Introduction
The panoramas represent a whole scene as an equirectangular map, delivering a 360 Field of View (FoV) of the surroundings. Hence, estimating depth from a single panoramic image can offer complete depth of the scene, with a holistic understanding of the surroundings. However, the equidistant unfolding characteristic (also known as panoramic distortion) of panoramas poses a new challenge for 360 depth estimation. Therefore, directly applying the perspective-based approaches [3, 4, 5, 6, 7, 8] to estimate depth from a single panorama is challenging. To address this issue, researchers [1, 9] have recently preferred to estimate accurate panoramic depth with the assistance of other projection formats (e.g., Cubemap projection (CP), and Tangent projection (TP)). These projections significantly reduce distortions by mapping a panorama onto several tangent faces, but damage the inherent continuous representation of panoramic images. Many approaches [1, 10, 11, 9] typically leverage the complementarity of these projection representations (shown in Fig. 1a) to estimate more accurate panoramic depth. Nevertheless, the trade-off between the FoV continuity and distortion elimination poses another challenge for the model design. Besides, these bi-projection fusion strategies require additional branches, increasing the computational burden. In view of this, we only take the equirectangular projection as the input and delve into the properties of the spherical space to deal with the distortion.
On the other hand, some approaches [12, 13, 2] deal with the panoramic distortions via modeling long-range dependencies (as illustrated in Fig. 1b). In previous work [12], long short-term memory [14] (LSTM) is introduced to disseminate global information in vertically compressed sequences. However, the geometric structure, which is an essential cue to estimate depth, is neglected, thus delivering blurred depth d. Meanwhile, those methods based on modeling long-range dependencies are either too global to perceive the local details [2], or too local to deal with the distortion efficiently [13].
@@
We evaluate the proposed SGFormer on two benchmarks, Structured3D [15] and Pano3D [16]. Experimental results demonstrate that our method outperforms existing methods by a large margin on both quantitative and qualitative results. Our main contributions can be summarized as follows:
@@ II-A Monocular 360 Depth Estimation
Bi-projection fusion strategy. BiFuse [1] introduces a dual-branch fusion method that combines CP and ERP features to alleviate distortion, while UniFuse [10] proposes a new framework that unidirectionally fuses CP features into ERP features during the decoding stage, enhancing both efficiency and performance. Glpanodepth [17] proposes a fusion module that combines ERP and CP features for panoramic depth estimation. 360MonoDepth [18] introduces a framework that uses tangent images for depth estimation from high-resolution 360 images. OmniFusion [11] directly converts 360 images into less-distorted perspective patches and uses geometric priors to merge the predictions of these patches into the ERP space to produce the final depth output in ERP format. HRDFuse [9] employs a collaborative mechanism that explores both the overall contextual information of the ERP and the local structure of the TP, estimating panoramic depth by learning the similarity of features between the ERP and TP.
Modeling long-range dependencies. SliceNet [12] divides the ERP image into vertical slices, however, it is hard to recover the details near the poles because it ignores the latitudinal distortion property. PanoFormer [13] attempts to estimate panoramic depth by leveraging a transformer-based architecture to process TP patches. By sampling the eight most relevant tokens for each central token on the ERP domain and adaptively adjusting local window sizes with a learnable offset, it enhances the precision in capturing the 2D structure and spatial details of each patch. EGFormer [2] leverages equirectangular geometry as the bias for local windows, thus estimating panoramic depth through an equirectangular geometry-aware local attention with a large receptive field. However, the vertical and horizontal shaped local windows also weaken the perception of the structure of each local area.
Other perspectives. Some researchers [19, 20, 21, 22, 23, 24, 25] employ various distortion-aware convolution filters to mitigate distortion, but due to the limited receptive fields, the final effect is still not entirely satisfactory. In deeper levels [26]. Several spherical CNNs [27] have employed convolution within the spectral domain, which could lead to a heavier computational burden. [28] considers the significance of the geometric structure in depth estimation and proposes a neural contourlet network for panoramic monocular depth estimation.
@@ II-B Transformer for Vision Tasks
Inspired by the transformer in the NLP field, the Vision Transformer [29] (ViT) uses a global self-attention mechanism for visual tasks. However, the global information interaction often results in higher computational costs. Swin Transformer [30] (SwinT) adopts a window-based self-attention, incorporating a hierarchical architecture and shifting window mechanism for image processing. CSwin Transformer [31] (CSwinT) introduces a cross-shaped window attention mechanism, conducting parallel self-attention computations on horizontal and vertical stripes. [32, 33] propose distortion relief methods specifically designed for panoramic image vision tasks. Despite these advancements, the distortion in panoramic images poses some challenges for transformers in panoramic image processing.
@@ III-A Architecture Overview
Figure 3: Overview of the query-based CGCPE module. In particular, the extracted features at the 3rd scale (denoted as f3) generate a global key (i.e. GCPE3) by passing through a transformer module with global spherical position embedding (GSPE). The global key is then processed through linear layers to generate keys corresponding to features at various resolutions. Features of different resolutions are used as queries, which are subsequently operated by $QK^{T}$. After another linear transformation, the corresponding GCPEs at different resolutions are obtained.
In Fig. 2, we show our complete framework motivated by the spherical properties. Overall, the proposed SGFormer takes a panorama as input and outputs a corresponding high-quality depth map. During the encoder stage, we employ Resnet-34 to extract features at five different resolutions (Fig. 2 ①). Then, these features are fed into the calculation of global conditional position embedding (CGCPE) module (Fig. 2 ② ), which is a query-based module that adaptively produces global conditional position embedding (GCPEs), compensating for additional spatial geometric structure (More details are presented in Section III-B). The visualization results of GCPEs demonstrate that they provide geometric details, promoting a comprehensive understanding of spatial location. Subsequently, a hierarchical decoder (the SPDecoder, described in Section III-C) is designed to estimate depth from these extracted features and integrates three spherical geometric priors to maintain the unity of the spherical geometric structure. As benefits by explicit structural cues and spherical priors (i.e. equidistortion, continuity, and surface distance), our scheme can effectively relieve distortion and these structure- and sphere-aware designs can promote the performance of our model.
@@ III-B Calculation of Global Conditional Position Embedding (CGCPE) Module
Specifically, as shown in Fig. 3, we select the feature with a $\frac{H}{16}\times\frac{W}{16}$ resolution (denoted as $f_{3}$) as the reference feature. There is a trade-off between computational efficiency and clear details. To balance them, we choose the middle scale. The selected features are fed into a transformer module, and then we achieve position embedding via a GSPE approach (more details are in section III-C, curve local embedding), thereby promoting the generation of panoramic structure (i.e., global key in Fig. 3). As shown in the top-left corner of Fig. 3, GSPE projects any two pixels in the ERP feature map into the spherical space to calculate the spherical surface distance as a sphere-aware embedding. Subsequently, the aggregated global geometric information is transformed into keys via a linear projection layer. At each scale, we leverage features as queries to query the keys, adaptively generating GCPE at each resolution. From Fig. 4, the depth details can be observed clearly, indicating that the query-based GCPE strategy can contribute to the spatial structure and depth details.
@@ III-C SPDecoder
Circular rotation (CR): The traditional transformer [34] architecture primarily relies on a global self-attention mechanism, which is inefficient when dealing with high-resolution images. To address this, SwinT [30] performs self-attention among divided window-based patches and employs a shifting window operation to facilitate feature interaction between the adjacent patches. Due to the unique seamless property of panoramic images, we implement self-attention interactions between local windows by rotating half the window size in the horizontal direction. Then we rotate back to the original state to align with the initial spatial representation. Since the north and south poles of panoramic images are not adjacent in space, we cannot directly apply CR in the vertical direction. However, benefiting from the BRP module, we can map the features from the vertical to the horizontal direction. Then, through CR, features that are originally in the vertical direction are interacted. Finally, we obtain a feature representation that combines horizontal and vertical directions.
@@
For each patch containing N points, we define $C$ as the CLE within each ${N\times N}$ patch. $\forall P_{1}\left(\lambda_{1},\varphi_{1}\right)$ and $P_{2}\left(\lambda_{2},\varphi_{2}\right)$ (i.e., the position of two points within a patch in spherical coordinates) to calculate the CLE. $c\left(P_{1},P_{2}\right)$ represents the distance between $P_{1}$ and $P_{2}$, where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi\in\left(-\pi,\pi\right)$ represents latitude/longitude, repectively. Through the Haversine formula [35], we can calculate $c\left(P_{1},P_{2}\right)$ as follows:
@@
where the radius (R) is set to 1. We employ the Eq. 5 and Eq. 6 to obtain the CLE within a patch.
@@
Dataset Method Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMS.lin $\downarrow$ RMSlog $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Bifuse [1] 0.0644 0.0565 0.4099 0.1194 0.9673 0.9892 0.9948
SliceNet [12] 0.1103 0.1273 0.6164 0.1811 0.9012 0.9705 0.9867
Yun et al. [36] 0.0505 0.0499 0.3475 0.1150 0.9700 0.9896 0.9947
S3D [15] PanoFormer [13] 0.0394 0.0346 0.2960 0.1004 0.9781 0.9918 0.9958
EGFormer [2] 0.0342 0.0279 0.2756 0.0932 0.9810 0.9928 0.9964
SGFormer (ours) 0.0303 0.0293 0.2429 0.0863 0.9857 0.9940 0.9969
Bifuse [1] 0.1704 0.1528 0.7272 0.2466 0.7680 0.9251 0.9731
SliceNet [12] 0.1254 0.1035 0.5761 0.1898 0.8575 0.9640 0.9867
Yun et al. [36] 0.0907 0.0658 0.4701 0.1502 0.9131 0.9792 0.9924
Pano3D [16] PanoFormer [13] 0.0699 0.0494 0.4046 0.1282 0.9436 0.9847 0.9939
EGFormer [2] 0.0660 0.0428 0.3874 0.1194 0.9503 0.9877 0.9952
SGFormer (ours) 0.0583 0.0364 0.3537 0.1101 0.9613 0.9896 0.9957
@@ IV-A Dataset and Implementation Details
Two datasets are employed in our evaluation: Structured3D [15] and Pano3D [16]. Structured3D is a high-quality synthetic dataset designed specifically to support a variety of structured 3D modeling tasks, providing many realistic images with detailed 3D structure annotations. On the other hand, the Pano3D dataset serves as a new benchmark for spherical panorama depth estimation. It aims to comprehensively evaluate the performance of depth prediction models, including the accuracy, the clarity of edges, and the smoothness of the image.
The reasons for not using other datasets are as follows: Firstly, the PanoSunCG dataset [38] is no longer publicly available. Secondly, the images in the 3D60 dataset [39] exhibit depth value leakage through their pixel brightness. Thirdly, Pano3D incorporates a rendered version of the Matterport3D [40], employing Pano3D and Matterport3D for assessment would result in unnecessary duplication. Lastly, in the Stanford2D3D dataset [41], the top and bottom parts of equirectangular images are not rendered properly. However, our method requires converting the north and south poles to the equatorial region through bipolar re-projection (BRP).
@@
Dataset Method Pub’Year Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMSE $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
EGFormer [2] ICCV’23 0.1528 0.1408 0.4974 0.8185 0.9338 0.9736
PanoFormer [13] ECCV’22 0.1122 0.0786 0.3945 0.8874 0.9584 0.9859
OmniFusion [11] CVPR’22 0.1154 0.0775 0.3809 0.8674 0.9603 0.9871
S2D3D [41] UniFuse [10] RAL’21 0.1124 0.0709 0.3555 0.8706 0.9704 0.9899
Elite360D [43] CVPR’24 0.1182 0.0728 0.3756 0.8872 0.9684 0.9892
SGFormer (ours) - 0.1040 0.0581 0.3406 0.8998 0.9693 0.9908
EGFormer [2] ICCV’23 0.1473 0.1517 0.6025 0.8158 0.9390 0.9735
PanoFormer [13] ECCV’22 0.1051 0.0966 0.4929 0.8908 0.9623 0.9831
BiFuse [1] CVPR’20 0.1126 0.0992 0.5027 0.8800 0.9613 0.9847
BiFuse++ [44] TPAMI’22 0.1123 0.0915 0.4853 0.8812 0.9656 0.9869
M3D [40] UniFuse [10] RAL’21 0.1144 0.0936 0.4835 0.8785 0.9659 0.9873
OmniFusion [11] CVPR’22 0.1161 0.1007 0.4931 0.8772 0.9615 0.9844
HRDFUSE [9] CVPR’23 0.1172 0.0971 0.5025 0.8674 0.9617 0.9849
Elite360D [43] CVPR’24 0.1115 0.0914 0.4875 0.8815 0.9646 0.9874
SGFormer (ours) - 0.1039 0.0865 0.4790 0.8946 0.9642 0.9859

### v3

@@ I Introduction
The panoramas represent a whole scene as an equirectangular map, delivering a 360 Field of View (FoV) of the surroundings. Hence, estimating depth from a single panoramic image can offer complete depth of the scene, with a holistic understanding of the surroundings. However, the equidistant unfolding characteristic (also known as panoramic distortion) of panoramas poses a new challenge for 360 depth estimation. Therefore, directly applying the perspective-based approaches [3, 4, 5, 6, 7, 8, 9] to estimate depth from a single panorama is challenging. To address this issue, researchers [1, 10] have recently preferred to estimate accurate panoramic depth with the assistance of other projection formats (e.g., Cubemap projection (CP), and Tangent projection (TP)). These projections significantly reduce distortions by mapping a panorama onto several tangent faces but damage the inherent continuous representation of panoramic images. Many approaches [1, 11, 12, 10] typically leverage the complementarity of these projection representations (shown in Fig. 1a) to estimate more accurate panoramic depth. Nevertheless, the trade-off between the FoV continuity and distortion elimination poses another challenge for the model design. Besides, these bi-projection fusion strategies require additional branches, increasing the computational burden. In view of this, we only take the equirectangular projection as the input and delve into the properties of the spherical space to deal with the distortion.
On the other hand, some approaches [13, 14, 2] deal with the panoramic distortions via modeling long-range dependencies (as illustrated in Fig. 1b). In previous work [13], long short-term memory [15] (LSTM) is introduced to disseminate global information in vertically compressed sequences. However, the geometric structure, which is an essential cue to estimate depth, is neglected, thus delivering blurred depth. Meanwhile, those methods based on modeling long-range dependencies are either too global to perceive the local details [2] or too local to deal with the distortion efficiently [14].
@@
We evaluate the proposed SGFormer on two benchmarks, Structured3D [16] and Pano3D [17]. Experimental results demonstrate that our method outperforms existing methods by a large margin on quantitative and qualitative results. Our main contributions can be summarized as follows:
@@ II-A Monocular 360 Depth Estimation
Bi-projection fusion strategy. BiFuse [1] introduces a dual-branch fusion method that combines CP and ERP features to alleviate distortion, while UniFuse [11] proposes a new framework that unidirectionally fuses CP features into ERP features during the decoding stage, enhancing both efficiency and performance. Glpanodepth [18] proposes a fusion module that combines ERP and CP features for panoramic depth estimation. 360MonoDepth [19] and [20] introduce the framework that uses tangent images for depth estimation from high-resolution 360 images. OmniFusion [12] directly converts 360 images into less-distorted perspective patches and uses geometric priors to merge the predictions of these patches into the ERP space to produce the final depth output in ERP format. HRDFuse [10] employs a collaborative mechanism that explores both the overall contextual information of the ERP and the local structure of the TP, estimating panoramic depth by learning the similarity of features between the ERP and TP.
Modeling long-range dependencies. SliceNet [13] divides the ERP image into vertical slices. However, it is hard to recover the details near the poles because it ignores the latitudinal distortion property. PanoFormer [14] attempts to estimate panoramic depth by leveraging a transformer-based architecture to process TP patches. By sampling the eight most relevant tokens for each central token on the ERP domain and adaptively adjusting local window sizes with a learnable offset, it enhances the precision in capturing the 2D structure and spatial details of each patch. EGFormer [2] leverages equirectangular geometry as the bias for local windows, thus estimating panoramic depth through an equirectangular geometry-aware local attention with a large receptive field. However, the vertical and horizontal shaped local windows also weaken the perception of the structure of each local area.
Other perspectives. Some researchers [21, 22, 23, 24, 25, 26, 27] employ various distortion-aware convolution filters to mitigate distortion, but due to the limited receptive fields, the final effect is still not entirely satisfactory. In deeper layers, several spherical CNNs [28, 29] have employed convolution within the spectral domain, which could lead to a heavier computational burden. [30] considers the significance of the geometric structure in depth estimation and proposes a neural contourlet network for panoramic monocular depth estimation. Different from supervised approaches, [31] proposes an unsupervised domain adaptation strategy that leverages unlabeled 360 data to fine-tune existing indoor panoramic depth estimation models (e.g., BiFuse++ or UniFuse). Specifically, it employs Depth Anything [32] to generate pseudo-labels for supervision. Hence, its performance is constrained by the base models and the quality of the pseudo-labels. In contrast, we integrate spherical geometric priors into the SPDecoder to design the base model. [33] introduces SegFuse, an end-to-end two-branch multi-task learning network designed for 360-depth estimation, which focuses more on outdoor than indoor scenarios. In driving scenarios, [34] proposes an innovative framework to address the challenges of imperfect observation and projection fusion in panoramic depth estimation. [35] presents the ADUULM-360 dataset, a multimodal dataset for depth estimation in adverse weather, which focuses more on outdoor driving scenarios. In addition, its dataset and code have not been released yet.
@@ II-B Transformer for Vision Tasks
Inspired by the transformer in the NLP field, the Vision Transformer [36] (ViT) uses a global self-attention mechanism for visual tasks. However, the global information interaction often results in higher computational costs. Swin Transformer [37] (SwinT) adopts a window-based self-attention, incorporating a hierarchical architecture and shifting window mechanism for image processing. [38] employs a convolution-free swin transformer as an image feature extractor. While SwinT is primarily used as a backbone in the architecture, our contribution is mainly focused on the decoder. CSwin Transformer [39] (CSwinT) introduces a cross-shaped window attention mechanism, conducting parallel self-attention computations on horizontal and vertical stripes. [40, 41] propose distortion relief methods specifically designed for panoramic image vision tasks. RCDFormer [42] integrates heterogeneous data from radar and cameras, leveraging a transformer architecture for dense depth estimation. Despite these advancements, the distortion in panoramic images poses some challenges for transformers in panoramic image processing. For example, the SwinT does not consider the distortion, equidistortion, and continuity of panoramic images. The SPDecoder is specially designed for panoramic images and can effectively cope with and utilize these unique properties.
@@
II-C 360 distortion processing and model lightweighting
To mitigate the adverse effects of distortion, [43] proposes a multi-projection fusion and refinement network. Recently, a simple and effective Cube Padding (CP) technique [44] has been proposed, demonstrating that the CP projection format can benefit panoramic tasks. For 360 video saliency detection, [45] introduces an innovative scheme based on the spherical convolutional neural network. DATFormer [46] addresses distortions with two modules: one for global distortion adaptation and another for reducing local distortions on multi-scale features. To handle distortions in spherical images, [47] presents a distortion-aware convolutional network, while [48] designs a spherical deep neural network (DNN) specifically to overcome distortions caused by the equirectangular projection. DINN [49] proposes a novel deformable invertible neural network (INN) for latitude-aware 360 image rescaling. New spherical image representations for 360 images are proposed by [50] and [51]. Additionally, [52] and [53] focus on developing lightweight models.
@@ III-A Architecture Overview
Figure 3: Overview of the query-based CGCPE module. In particular, the extracted feature f3 generates a global key (i.e. GCPE3) by passing through a transformer module with global spherical position embedding (GSPE). The global key is then processed through linear layers to generate keys corresponding to features at various resolutions. The original features of different resolutions are used as queries, which are subsequently operated by $QK^{T}$. After another linear transformation, the corresponding GCPEs at different resolutions are obtained.
In Fig. 2, we show our complete framework motivated by the spherical properties. Overall, the proposed SGFormer takes a panorama as input and outputs a corresponding high-quality depth map. During the encoder stage, we employ Resnet-34 to extract features at five different resolutions (Fig. 2 ①). Then, these features are fed into the calculation of the global conditional position embedding (CGCPE) module (Fig. 2 ② ), which is a query-based module that adaptively produces global conditional position embedding (GCPEs), compensating for additional spatial geometric structure (More details are presented in Section III-B). The visualization results of GCPEs demonstrate that they provide geometric details, promoting a comprehensive understanding of spatial location. Subsequently, a hierarchical decoder (the SPDecoder, described in Section III-C) is designed to estimate depth from these extracted features and integrates three spherical geometric priors to maintain the unity of the spherical geometric structure. As benefits by explicit structural cues and spherical priors (i.e. equidistortion, continuity, and surface distance), our scheme can effectively relieve distortion, and these structure- and sphere-aware designs can promote the performance of our model.
@@ III-B Calculation of Global Conditional Position Embedding (CGCPE) Module
Specifically, as shown in Fig. 3, we select the feature with a $\frac{H}{16}\times\frac{W}{16}$ resolution (denoted as $f_{3}$) as the reference feature. There is a trade-off between computational efficiency and clear details. To balance them, we choose the middle scale. The selected features are fed into a transformer module. Then we achieve position embedding via a GSPE approach (more details are in section III-C, curve local embedding), thereby promoting the generation of panoramic structure (i.e., global key in Fig. 3). As shown in the top-left corner of Fig. 3, GSPE projects any two pixels in the ERP feature map into the spherical space to calculate the spherical surface distance as a sphere-aware embedding. Subsequently, the aggregated global geometric information is transformed into keys via a linear projection layer. At each scale, we leverage features as queries to query the keys, adaptively generating GCPE at each resolution.
@@ III-C SPDecoder
Circular rotation (CR): The traditional transformer [54] architecture primarily relies on a global self-attention mechanism, which is inefficient when dealing with high-resolution images. To address this, SwinT [37] performs self-attention among divided window-based patches and employs a shifting window operation to facilitate feature interaction between the adjacent patches. Due to the unique seamless property of panoramic images, we implement self-attention interactions between local windows by rotating half the window size in the horizontal direction. Then, we rotate back to the original state to align with the initial spatial representation. Since the north and south poles of panoramic images are not adjacent in space, we cannot directly apply CR in the vertical direction. However, benefiting from the BRP module, we can map the features from the vertical to the horizontal direction. Then, through CR, features that are originally in the vertical direction are interacted. Finally, we obtain a feature representation that combines horizontal and vertical directions.
@@
For each patch containing N points, we define $C$ as the CLE within each ${N\times N}$ patch. $\forall P_{1}\left(\lambda_{1},\varphi_{1}\right)$ and $P_{2}\left(\lambda_{2},\varphi_{2}\right)$ (i.e., the position of two points within a patch in spherical coordinates) to calculate the CLE. $c\left(P_{1},P_{2}\right)$ represents the distance between $P_{1}$ and $P_{2}$, where $\lambda\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$/$\varphi\in\left(-\pi,\pi\right)$ represents latitude/longitude, respectively. Through the Haversine formula [55], we can calculate $c\left(P_{1},P_{2}\right)$ as follows:
@@
where the radius (R) is set to 1. We employ Eq. 5 and Eq. 6 to obtain the CLE within a patch.
@@
Dataset Method Pub’Year Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMS.lin $\downarrow$ RMSlog $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
Bifuse [1] CVPR’20 0.0644 0.0565 0.4099 0.1194 0.9673 0.9892 0.9948
SliceNet [13] CVPR’21 0.1103 0.1273 0.6164 0.1811 0.9012 0.9705 0.9867
Yun et al. [56] AAAI’22 0.0505 0.0499 0.3475 0.1150 0.9700 0.9896 0.9947
S3D [16] PanoFormer [14] ECCV’22 0.0394 0.0346 0.2960 0.1004 0.9781 0.9918 0.9958
EGFormer [2] ICCV’23 0.0342 0.0279 0.2756 0.0932 0.9810 0.9928 0.9964
Elite360D [57] CVPR’24 0.0391 0.0350 0.2804 0.0971 0.9799 0.9925 0.9962
SGFormer (ours) - 0.0303 0.0293 0.2429 0.0863 0.9857 0.9940 0.9969
Bifuse [1] CVPR’20 0.1704 0.1528 0.7272 0.2466 0.7680 0.9251 0.9731
SliceNet [13] CVPR’21 0.1254 0.1035 0.5761 0.1898 0.8575 0.9640 0.9867
Yun et al. [56] AAAI’22 0.0907 0.0658 0.4701 0.1502 0.9131 0.9792 0.9924
Pano3D [17] PanoFormer [14] ECCV’22 0.0699 0.0494 0.4046 0.1282 0.9436 0.9847 0.9939
EGFormer [2] ICCV’23 0.0660 0.0428 0.3874 0.1194 0.9503 0.9877 0.9952
Elite360D [57] CVPR’24 0.0713 0.0438 0.3888 0.1236 0.9482 0.9870 0.9950
SGFormer (ours) - 0.0583 0.0364 0.3537 0.1101 0.9613 0.9896 0.9957
@@ IV-A Dataset and Implementation Details
Two datasets are employed in our evaluation: Structured3D [16] and Pano3D [17]. Structured3D is a high-quality synthetic dataset designed specifically to support a variety of structured 3D modeling tasks, providing many realistic images with detailed 3D structure annotations. On the other hand, the Pano3D dataset serves as a new benchmark for spherical panorama depth estimation. It aims to comprehensively evaluate the performance of depth prediction models, including the accuracy, the clarity of edges, and the image’s smoothness. Consistent with EGFormer [2], we follow the official dataset split [16, 17]. For the Structured3D dataset, we use scene_00000 to scene_02999 for training, scene_03000 to scene_03249 for validation, and scene_3250 to scene_03499 for testing. For the Pano3D dataset, we use 5636 images for training, 744 images for validation, and 1527 images for testing.
The reasons for not using PanoSunCG and 3D60 datasets are as follows: Firstly, the PanoSunCG dataset [59] is no longer publicly available. Secondly, the images in the 3D60 dataset [60] exhibit depth value leakage through their pixel brightness.
@@
Dataset Method Pub’Year Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
EGFormer [2] ICCV’23 0.1528 0.1408 0.4974 0.2803 0.8185 0.9338 0.9736
PanoFormer [14] ECCV’22 0.1122 0.0786 0.3945 0.2119 0.8874 0.9584 0.9859
OmniFusion [12] CVPR’22 0.1154 0.0775 0.3809 0.2164 0.8674 0.9603 0.9871
S2D3D [62] UniFuse [11] RAL’21 0.1124 0.0709 0.3555 0.1978 0.8706 0.9704 0.9899
Elite360D [57] CVPR’24 0.1182 0.0728 0.3756 0.2219 0.8872 0.9684 0.9892
SGFormer (ours) - 0.1040 0.0581 0.3406 0.2017 0.8998 0.9693 0.9908
EGFormer [2] ICCV’23 0.1473 0.1517 0.6025 0.3589 0.8158 0.9390 0.9735
PanoFormer [14] ECCV’22 0.1051 0.0966 0.4929 0.2777 0.8908 0.9623 0.9831
BiFuse [1] CVPR’20 0.1126 0.0992 0.5027 0.2918 0.8800 0.9613 0.9847
BiFuse++ [63] TPAMI’22 0.1123 0.0915 0.4853 0.2783 0.8812 0.9656 0.9869
M3D [64] UniFuse [11] RAL’21 0.1144 0.0936 0.4835 0.2821 0.8785 0.9659 0.9873
OmniFusion [12] CVPR’22 0.1161 0.1007 0.4931 0.2870 0.8772 0.9615 0.9844
HRDFUSE [10] CVPR’23 0.1172 0.0971 0.5025 0.2963 0.8674 0.9617 0.9849
Elite360D [57] CVPR’24 0.1115 0.0914 0.4875 0.2826 0.8815 0.9646 0.9874
SGFormer (ours) - 0.1039 0.0865 0.4790 0.2748 0.8946 0.9642 0.9859